# Classification of abelian Schur groups I

### v1

@@ 7.3. Dense $S$-rings
Therefore $m\leq 2$. If $m=1$, then $U$ is trivial, a contradiction to the assumption of the proposition. If $m=2$, then $U=L=P$ is an $\mathcal{A}$-subgroup. By Lemma 2.1 and Eq. (28), we obtain $P=L\leq\operatorname{rad}(X)$, a contradiction to the assumption of the claim. ∎
If $L\leq\operatorname{rad}(X)$ for every $X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\times U)}$, then $\mathcal{A}$ is the nontrivial $(H\times U)/L$-wreath product and Statement $(1)$ of the proposition holds.
Suppose that $L\nleq\operatorname{rad}(X)$ for some $X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\times U)}$. By Claim $4$, we have $n=3^{k}$, $\mathcal{A}_{H}=\mathcal{T}_{H}$, and $L=P$. Let $W$ be the $\mathcal{A}_{D}$-subgroup generated by $X_{D}$, where $X$ runs over all basic sets of $\mathcal{A}$ outside $H\times U$ such that $P\nleq\operatorname{rad}(X)$. Note that $W$ is indeed an $\mathcal{A}_{D}$-subgroup by Lemma 3.2(1). By the definition, $W>U$. Clearly, $\mathcal{A}$ is the $(H\times W)/P$-wreath product. One can see that $\operatorname{rad}(\mathcal{A}_{W})=P$ by Claim $4$.
In view of the above paragraph, to prove Statement $(2)$ of the proposition, it is enough to show that $\mathcal{A}_{H\times W}$ is cyclotomic. Let $X\in\mathcal{S}(\mathcal{A}_{H\times W})_{G\setminus(H\times U)}$ such that $P\nleq\operatorname{rad}(X)$ and $X_{D}$ contains a generator of $W$. By Claim $4$, the set $X$ is one of the forms from Eq. (27), where $x\in W$ is a generator of $W$. Let $\sigma\in\operatorname{Aut}(H)$ such that $\sigma=(h_{1}h_{2}h_{3})$ and $\tau_{1},\tau_{2}\in\operatorname{Aut}(W)$ such that $x^{\tau_{1}}=xx_{0}$ and $x^{\tau_{2}}=x^{-1}$. Put
@@
where $\varphi,\psi\in\operatorname{Aut}(H\times W)$ are such that $\varphi^{H}=\sigma$, $\varphi^{D}=\tau_{1}$, $\psi^{H}=\operatorname{id}_{H}$, $\psi^{D}=\tau_{2}$. Clearly, $K_{1}\cong C_{3}$ and $K_{2}\cong C_{6}$. Due to the definitions of $K_{1}$ and $K_{2}$, we conclude that $X\in\operatorname{Orb}(K,H\times W)$ and $|X|=|K|$, where $K=K_{1}$ if $X$ is of the first form from Eq. (27) and $K=K_{2}$ otherwise. It is easy to see that $H^{\#},X_{D}\in\operatorname{Orb}(K,H\times W)$. Put $X^{0}=X$ and $X^{i}=(X^{i-1})^{[p]}$ for $i\in\{1,\ldots,l-1\}$, where $l$ is such that $|W|=3^{l}$. Eq. (27) implies that $X^{i}=H^{\#}x^{p^{i}}$ or $X^{i}=H^{\#}x^{p^{i}}\cup H^{\#}x^{-p^{i}}$ for $i\in\{1,\ldots,l-1\}$. Each $X^{i}$ is an $\mathcal{A}$-set by Lemma 2.4. Moreover, using Lemma 2.1 and Lemma 3.2(1), it is easy to verify that each $X^{i}$ and $X^{i}_{D}$ are exactly basic sets of $\mathcal{A}_{H\times W}$ and $\mathcal{A}_{W}$, respectively. Due to the definition of $K$, each $X^{i}$ and $X^{i}_{D}$ are orbits of $K$. Every $Y\in\mathcal{S}(\mathcal{A}_{H\times W})_{G\setminus H}$ is rationally conjugate to $X^{i}$ or $X^{i}_{D}$ for some $i\in\{0,\ldots,l-1\}$ by Lemma 2.3 and hence $Y\in\operatorname{Orb}(K,H\times W)$. Thus, every basic set of $\mathcal{A}_{H\times W}$ is an orbit of $K$, i.e. $\mathcal{A}_{H\times W}=\operatorname{Cyc}(K,G)$ as required. ∎
@@
One can see that $\operatorname{Aut}(\mathcal{A})^{H}\approx_{2}\operatorname{Aut}(\mathcal{A}_{H})\cong\operatorname{Sym}(4)$ because $\mathcal{A}$ is schurian. Due to Remark 2.7, we conclude that $|\operatorname{Aut}(\mathcal{A})^{H}|\in\{12,24\}$.
Since $\mathcal{A}\neq\mathcal{A}_{H}\otimes\mathcal{A}_{D}$, there is $X\in\mathcal{S}(\mathcal{A})_{G\setminus(U\cup D)}$ such that $X\neq X_{H}\times X_{D}$. Lemma 7.10 implies that $\lambda_{X}=1$. From Lemma 3.2(1) it follows that $X_{D}\in\mathcal{S}(\mathcal{A}_{D})$ and $X_{H}=H^{\#}$. Let $L=\operatorname{rad}(X)$, $U=\langle X\rangle$, and $W=\langle X_{D}\rangle$. The assumptions of the lemma yield that $L\leq D$. The image of $T\subseteq G$ under the canonical epimorphism from $G$ to $G/L$ is denoted by $\mkern 1.5mu\overline{\mkern-1.5muT\mkern-1.5mu}\mkern 1.5mu$. The argument from the previous paragraph applied to $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu}$ implies that $|\operatorname{Aut}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu})^{\mkern 1.5mu\overline{\mkern-1.5muH\mkern-1.5mu}\mkern 1.5mu}|\in\{12,24\}$. To prove the lemma, it is enough to show that
@@
$|\operatorname{Aut}(\mathcal{A})^{H}|=|\operatorname{Aut}(\mathcal{A})^{\mkern 1.5mu\overline{\mkern-1.5muH\mkern-1.5mu}\mkern 1.5mu}|\leq|\operatorname{Aut}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu})^{\mkern 1.5mu\overline{\mkern-1.5muH\mkern-1.5mu}\mkern 1.5mu}|=12$
as required, where the first equality holds because $L\leq D$.
@@
As $\lambda_{X}=1$ and $L\leq D$, we obtain $\lambda_{\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu}=|\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\cap\mkern 1.5mu\overline{\mkern-1.5muH\mkern-1.5mu}\mkern 1.5mu\mkern 1.5mu\overline{\mkern-1.5mux\mkern-1.5mu}\mkern 1.5mu|=1$ for every $\mkern 1.5mu\overline{\mkern-1.5mux\mkern-1.5mu}\mkern 1.5mu\in\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu$. So $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu}\neq\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muH\mkern-1.5mu}\mkern 1.5mu}\otimes\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muW\mkern-1.5mu}\mkern 1.5mu}$. The set $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{D}$ is a basic set of $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muD\mkern-1.5mu}\mkern 1.5mu}$. If $|\operatorname{rad}(\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{D})|=1$, then $|\operatorname{rad}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muW\mkern-1.5mu}\mkern 1.5mu})|=1$ and Proposition 7.11(2) applied to $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu}$ yields that $|\operatorname{Aut}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu})^{\mkern 1.5mu\overline{\mkern-1.5muH\mkern-1.5mu}\mkern 1.5mu}|=12$ as desired.
Suppose that $|\operatorname{rad}(\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{D})|>1$. In view of Lemma 6.12 if $|W|=pq$ and Lemma 6.10 if $|W|$ is a $p$-power, we conclude that Proposition 7.12 holds for $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu}$. If Statement $(1)$ of Proposition 7.12 holds for $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu}$, then $|\operatorname{rad}(\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu)|>1$, a contradiction to Eq. (29). Therefore Statement $(2)$ of Proposition 7.12 holds for $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu}$. Due to Claim $4$ and Eq. (29), the basic sets of $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu}$ inside $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{H}\times\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{D}$ are of the form
@@
where the $\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{i}$’s are pairwise distinct elements of $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{H}$ and the $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{i}$’s are pairwise disjoint subsets of $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{D}$ of size at most $2$. If $|\operatorname{Aut}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu})^{\mkern 1.5mu\overline{\mkern-1.5muH\mkern-1.5mu}\mkern 1.5mu}|=24$, then there is $f\in\operatorname{Aut}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu})$ such that $\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{1}^{f}=\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{1}$, $\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{2}^{f}=\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{3}$, and $\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{3}^{f}=\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{2}$. Eq. (2) implies that
@@
If $H$ ($D$, respectively) is not an $\mathcal{A}$-subgroup, then Proposition 7.3 and Proposition 7.5 (Proposition 7.4 and Proposition 7.7, respectively) imply that Theorem 7.1 holds for $\mathcal{A}$ or $\mathcal{A}$ is a nontrivial $S$-wreath product for an $\mathcal{A}$-section $S=U/L$ such that $4\nmid|W|$ and $|\operatorname{rad}(\mathcal{A}_{W})|=1$, where $W=U$ if $H$ is not an $\mathcal{A}$-subgroup and $W=G/L$ if $D$ is not an $\mathcal{A}$-subgroup. Note that $|W_{H}|\leq 2$, where $W_{H}$ is the Sylow $2$-subgroup of $W$, because $|W|$ is not divisible by $4$ and consequently $\mathcal{A}_{W_{H}}=\mathbb{Z}W_{H}$. So Lemma 6.2 and Lemma 3.2(2) yield that
@@
where $W_{D}$ is the Hall $2^{\prime}$-subgroup of $W$ (here $W_{H}$ and $W_{D}$ can be trivial). Since $|\operatorname{rad}(\mathcal{A}_{W})|=1$, we conclude that $|\operatorname{rad}(\mathcal{A}_{W_{D}})|=1$. If $\mathcal{A}_{W_{D}}=\mathcal{T}_{W_{D}}$, then due to Eq. (30), we obtain $L\in\{W_{H},W_{D},W\}$ whenever $W=U$ and $S\in\{W_{H},W_{D},\{L\}\}$ whenever $W=G/L$. This yields that Statement $(1)$ or $(2)$ from Theorem 7.1 holds. Suppose that $\mathcal{A}_{W_{D}}\neq\mathcal{T}_{W_{D}}$. Then $\mathcal{A}_{W_{D}}=\operatorname{Cyc}(K_{D},W_{D})$ for some $K_{D}\leq\operatorname{Aut}(W_{D})$ by Lemma 6.2. As $|W_{H}|\leq 2$, we have $\mathcal{A}_{W_{H}}=\operatorname{Cyc}(K_{H},W_{H})$, where $K_{H}$ is trivial. Therefore $\mathcal{A}_{W}=\operatorname{Cyc}(K_{H}\times K_{D},W)$ by Eqs. (4) and (30) and Statement $(3)$ from Theorem 7.1 holds.
Suppose that $H$ and $D$ are $\mathcal{A}$-subgroups. If $|\operatorname{rad}(\mathcal{A}_{D})|=1$, then $\mathcal{A}=\mathcal{A}_{H}\otimes\mathcal{A}_{D}$ or $\mathcal{A}$ is cyclotomic by Proposition 7.11(1) and we are done. If $|\operatorname{rad}(\mathcal{A}_{D})|>1$, then $\mathcal{A}_{D}$ is a nontrivial $U/L$-wreath product for some $\mathcal{A}_{D}$-section $U/L$ such that $L$ is a minimal nontrivial $\mathcal{A}_{D}$-subgroup and $|\operatorname{rad}(\mathcal{A}_{U})|=1$. Indeed, this follows from Lemma 6.12 if $n=pq$ and from Lemma 6.10 if $n=p^{k}$. So Proposition 7.12 holds for $\mathcal{A}$. If Statement $(2)$ of Proposition 7.12 holds for $\mathcal{A}$, then $n=3^{k}$, $\mathcal{A}$ is the $S=(H\times W)/L$-wreath product (possibly, trivial) for an $\mathcal{A}_{D}$-subgroup $W>U$ such that $|\operatorname{rad}(\mathcal{A}_{W})|=3$, and $\mathcal{A}_{H\times W}$ is cyclotomic. In this case, $H\times W=G$ and $\mathcal{A}$ is cyclotomic or $H\times W<G$, $\mathcal{A}$ is the nontrivial $S$-wreath product, and Statement $(4)$ from Theorem 7.1 holds as required.
Suppose that Statement $(1)$ of Proposition 7.12 holds for $\mathcal{A}$, i.e. $\mathcal{A}$ is the nontrivial $S=V/L$-wreath product, where $V=H\times U$. If $\mathcal{A}_{V}\neq\mathcal{A}_{H}\otimes\mathcal{A}_{U}$, then $\mathcal{A}_{V}$ is cyclotomic by Proposition 7.11(1). So Statement $(4)$ from Theorem 7.1 holds for $\mathcal{A}$ and $S$. Further, we assume that

### v3

@@ 7.3. Dense $S$-rings
Therefore $m\leq 2$. If $m=1$, then $U$ is trivial. However, this contradicts to the assumption of the theorem that $\mathcal{A}$ is the nontrivial $U/L$-wreath product. If $m=2$, then $U=L=P$ is an $\mathcal{A}$-subgroup. By Lemma 2.1 and Eq. (32), we obtain $P=L\leq\operatorname{rad}(X)$, a contradiction to the assumption of the claim. ∎
If $L\leq\operatorname{rad}(X)$ for every $X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\times U)}$, then $\mathcal{A}$ is the nontrivial $(H\times U)/L$-wreath product and Statement $(1)$ of the theorem holds.
Suppose that $L\nleq\operatorname{rad}(X)$ for some $X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\times U)}$. As $L$ is a minimal nontrivial $\mathcal{A}_{D}$-subgroup, we obtain $P\nleq\operatorname{rad}(X)$. So the set
$\mathcal{W}=\{X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\times U)}:~P\nleq\operatorname{rad}(X)\}$
is nonempty. Put
$W=\langle X_{D}:~X\in\mathcal{W}\rangle.$
Observe that $X_{D}$ is an $\mathcal{A}_{D}$-set for every $X\in\mathcal{W}$ by Lemma 3.2(1) and hence $W$ is an $\mathcal{A}_{D}$-subgroup. By the definition, $W>U$. Claim $4$ implies that $n=3^{k}$, $\mathcal{A}_{H}=\mathcal{T}_{H}$, $L=P$, and $|X_{D}|\leq 6$ for every $X\in\mathcal{W}$. Together with $\operatorname{rad}(X_{D})\geq L$, the latter yields that $\operatorname{rad}(X_{D})=P$ for every $X\in\mathcal{W}$ and consequently $\operatorname{rad}(\mathcal{A}_{W})=P$. By the definition of $W$, the radical of every basic set outside $H\times W$ contains $P$. Therefore $\mathcal{A}$ is the $(H\times W)/P$-wreath product (possibly, trivial).
In view of the above paragraph, to prove Statement $(2)$ of the theorem, it suffices to show that the $S$-ring $\mathcal{A}_{H\times W}$ is cyclotomic. Let $X\in\mathcal{W}$ (recall that $\mathcal{W}$ is nonempty) be such that $X_{D}$ contains a generator of $W$. By Claim $4$, the set $X$ is one of the forms from Eq. (31), where $x\in W$ is a generator of $W$. Let $\sigma\in\operatorname{Aut}(H)$ be such that $\sigma=(h_{1}h_{2}h_{3})$, and let $\tau_{1},\tau_{2}\in\operatorname{Aut}(W)$ be such that $x^{\tau_{1}}=xx_{0}$ and $x^{\tau_{2}}=x^{-1}$. Put
@@
where $\varphi,\psi\in\operatorname{Aut}(H\times W)$ are such that $\varphi^{H}=\sigma$, $\varphi^{D}=\tau_{1}$, $\psi^{H}=\operatorname{id}_{H}$, $\psi^{D}=\tau_{2}$. Clearly, $K_{1}\cong C_{3}$ and $K_{2}\cong C_{6}$. By the definitions of $K_{1}$ and $K_{2}$, we have $X\in\operatorname{Orb}(K,H\times W)$ and $|X|=|K|$, where $K=K_{1}$ if $X$ is of the first form from Eq. (31) and $K=K_{2}$ otherwise. It is easy to see that $H^{\#},X_{D}\in\operatorname{Orb}(K,H\times W)$. Put $X^{0}=X$ and $X^{i}=(X^{i-1})^{[p]}$ for $i\in\{1,\ldots,l-1\}$, where $l$ is such that $|W|=3^{l}$. Eq. (31) implies that $X^{i}=H^{\#}x^{p^{i}}$ or $X^{i}=H^{\#}x^{p^{i}}\cup H^{\#}x^{-p^{i}}$ for $i\in\{1,\ldots,l-1\}$. Each $X^{i}$ is an $\mathcal{A}$-set by Lemma 2.4. Moreover, using Lemma 2.1 and Lemma 3.2(1), it is easy to verify that each $X^{i}$ and $X^{i}_{D}$ are exactly basic sets of $\mathcal{A}_{H\times W}$ and $\mathcal{A}_{W}$, respectively. Due to the definition of $K$, each $X^{i}$ and $X^{i}_{D}$ are orbits of $K$. Every $Y\in\mathcal{S}(\mathcal{A}_{H\times W})_{G\setminus H}$ is rationally conjugate to $X^{i}$ or $X^{i}_{D}$ for some $i\in\{0,\ldots,l-1\}$ by Lemma 2.3 and hence $Y\in\operatorname{Orb}(K,H\times W)$. Thus, every basic set of $\mathcal{A}_{H\times W}$ is an orbit of $K$, i.e. $\mathcal{A}_{H\times W}=\operatorname{Cyc}(K,G)$ as required. ∎
@@
One can see that $\operatorname{Aut}(\mathcal{A})^{H}\approx_{2}\operatorname{Aut}(\mathcal{A}_{H})\cong\operatorname{Sym}(4)$ because $\mathcal{A}$ is schurian. Moreover, $|\operatorname{Aut}(\mathcal{A})^{H}|\in\{12,24\}$ by Remark 2.7.
Since $\mathcal{A}\neq\mathcal{A}_{H}\otimes\mathcal{A}_{D}$, there is $X\in\mathcal{S}(\mathcal{A})_{G\setminus(U\cup D)}$ such that $X\neq X_{H}\times X_{D}$. Lemma 7.11 implies that $\lambda_{X}=1$. From Lemma 3.2(1) it follows that $X_{D}\in\mathcal{S}(\mathcal{A}_{D})$ and $X_{H}=H^{\#}$. Let $L=\operatorname{rad}(X)$, $U=\langle X\rangle$, and $W=\langle X_{D}\rangle$. Since $D$ is an $\mathcal{A}$-subgroup, we have $L\leq D$. The image of $T\subseteq G$ under the canonical epimorphism from $G$ to $G/L$ is denoted by $\mkern 1.5mu\overline{\mkern-1.5muT\mkern-1.5mu}\mkern 1.5mu$. The argument from the previous paragraph applied to $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu}$ implies that $|\operatorname{Aut}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu})^{\mkern 1.5mu\overline{\mkern-1.5muH\mkern-1.5mu}\mkern 1.5mu}|\in\{12,24\}$. To prove the lemma, it is enough to show that
@@
$|\operatorname{Aut}(\mathcal{A})^{H}|=|\operatorname{Aut}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu})^{\mkern 1.5mu\overline{\mkern-1.5muH\mkern-1.5mu}\mkern 1.5mu}|\leq|\operatorname{Aut}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu})^{\mkern 1.5mu\overline{\mkern-1.5muH\mkern-1.5mu}\mkern 1.5mu}|=12$
as required, where the first equality holds because $|L\cap H|=1$ and hence the mapping $f\mapsto\mkern 1.5mu\overline{\mkern-1.5muf\mkern-1.5mu}\mkern 1.5mu$ from $\operatorname{Aut}(\mathcal{A})^{H}$ to $\operatorname{Aut}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu})^{\mkern 1.5mu\overline{\mkern-1.5muH\mkern-1.5mu}\mkern 1.5mu}$, where $\mkern 1.5mu\overline{\mkern-1.5muf\mkern-1.5mu}\mkern 1.5mu$ is the permutation induced by $f$ on $\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu$, is an isomorphism.
@@
As $\lambda_{X}=1$ and $L\leq D$, we obtain $\lambda_{\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu}=|\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\cap\mkern 1.5mu\overline{\mkern-1.5muH\mkern-1.5mu}\mkern 1.5mu\mkern 1.5mu\overline{\mkern-1.5mux\mkern-1.5mu}\mkern 1.5mu|=1$ for every $\mkern 1.5mu\overline{\mkern-1.5mux\mkern-1.5mu}\mkern 1.5mu\in\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu$. So $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu}\neq\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muH\mkern-1.5mu}\mkern 1.5mu}\otimes\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muW\mkern-1.5mu}\mkern 1.5mu}$. The set $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{D}$ is a basic set of $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muD\mkern-1.5mu}\mkern 1.5mu}$. If $|\operatorname{rad}(\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{D})|=1$, then $|\operatorname{rad}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muW\mkern-1.5mu}\mkern 1.5mu})|=1$ and Theorem 7.12(2) applied to $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu}$ yields that $|\operatorname{Aut}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu})^{\mkern 1.5mu\overline{\mkern-1.5muH\mkern-1.5mu}\mkern 1.5mu}|=12$ as desired.
Suppose that $|\operatorname{rad}(\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{D})|>1$. In view of Lemma 6.12 if $|W|=pq$ and Lemma 6.10 if $|W|$ is a $p$-power, Theorem 7.13 holds for $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu}$. If Statement $(1)$ of Theorem 7.13 holds for $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu}$, then $|\operatorname{rad}(\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu)|>1$, a contradiction to Eq. (33). Therefore Statement $(2)$ of Theorem 7.13 holds for $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu}$. Due to Claim $4$ and Eq. (33), the basic sets of $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu}$ inside $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{H}\times\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{D}$ are of the form
@@
where the $\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{i}$ are pairwise distinct elements of $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{H}$ and the $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{i}$ are pairwise disjoint subsets of $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{D}$ of size at most $2$. If $|\operatorname{Aut}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu})^{\mkern 1.5mu\overline{\mkern-1.5muH\mkern-1.5mu}\mkern 1.5mu}|=24$, then there is $f\in\operatorname{Aut}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu})$ such that $\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{1}^{f}=\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{1}$, $\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{2}^{f}=\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{3}$, and $\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{3}^{f}=\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{2}$. Eq. (2) implies that
@@
Suppose that $H$ or $D$ is not an $\mathcal{A}$-subgroup. Then Theorem 7.1 holds for $\mathcal{A}$ or $\mathcal{A}$ is a nontrivial $S$-wreath product for an $\mathcal{A}$-section $S=U/L$ such that $4\nmid|W|$ and $|\operatorname{rad}(\mathcal{A}_{W})|=1$, where $W=U$ if $H$ is not an $\mathcal{A}$-subgroup and $W=G/L$ if $D$ is not an $\mathcal{A}$-subgroup: indeed, this follows from Theorem 7.3 in the former case and from Theorem 7.8 in the latter one.
Note that $|W_{H}|\leq 2$, where $W_{H}$ is the Sylow $2$-subgroup of $W$, because $|W|$ is not divisible by $4$ and consequently $\mathcal{A}_{W_{H}}=\mathbb{Z}W_{H}$. So Lemma 3.2(2) yields that
@@
where $W_{D}$ is the Hall $2^{\prime}$-subgroup of $W$ (here $W_{H}$ and $W_{D}$ can be trivial). Since $|\operatorname{rad}(\mathcal{A}_{W})|=1$, we conclude that $|\operatorname{rad}(\mathcal{A}_{W_{D}})|=1$.
If $\mathcal{A}_{W_{D}}=\mathcal{T}_{W_{D}}$, then due to Eq. (34), we obtain $L\in\{W_{H},W_{D},W\}$ whenever $W=U$ and $S\in\{W_{H},W_{D},\{L\}\}$ whenever $W=G/L$. This yields that Statement $(1)$ or $(2)$ from Theorem 7.1 holds. If $\mathcal{A}_{W_{D}}\neq\mathcal{T}_{W_{D}}$, then $\mathcal{A}_{W_{D}}=\operatorname{Cyc}(K_{D},W_{D})$ for some $K_{D}\leq\operatorname{Aut}(W_{D})$ by Lemma 6.2. As $|W_{H}|\leq 2$, we have $\mathcal{A}_{W_{H}}=\operatorname{Cyc}(K_{H},W_{H})$, where $K_{H}$ is trivial. Therefore $\mathcal{A}_{W}=\operatorname{Cyc}(K_{H}\times K_{D},W)$ by Eqs. (4) and (34) and Statement $(3)$ from Theorem 7.1 holds.
Now suppose that $H$ and $D$ are $\mathcal{A}$-subgroups. If $|\operatorname{rad}(\mathcal{A}_{D})|=1$, then $\mathcal{A}=\mathcal{A}_{H}\otimes\mathcal{A}_{D}$ or $\mathcal{A}$ is cyclotomic by Theorem 7.12(1) and we are done. If $|\operatorname{rad}(\mathcal{A}_{D})|>1$, then $\mathcal{A}_{D}$ is a nontrivial $U/L$-wreath product for some $\mathcal{A}_{D}$-section $U/L$ such that $L$ is a minimal nontrivial $\mathcal{A}_{D}$-subgroup and $|\operatorname{rad}(\mathcal{A}_{U})|=1$. Indeed, this follows from Lemma 6.12 if $n=pq$ and from Lemma 6.10 if $n=p^{k}$. So Theorem 7.13 holds for $\mathcal{A}$. If Statement $(2)$ of Theorem 7.13 holds for $\mathcal{A}$, then $n=3^{k}$, $\mathcal{A}$ is the $S=(H\times W)/L$-wreath product (possibly, trivial) for an $\mathcal{A}_{D}$-subgroup $W>U$ such that $|\operatorname{rad}(\mathcal{A}_{W})|=3$, and $\mathcal{A}_{H\times W}$ is cyclotomic. In this case, $H\times W=G$ and $\mathcal{A}$ is cyclotomic or $H\times W<G$, $\mathcal{A}$ is the nontrivial $S$-wreath product. Thus, Statement $(4)$ from Theorem 7.1 holds.
Suppose that Statement $(1)$ of Theorem 7.13 holds for $\mathcal{A}$, i.e. $\mathcal{A}$ is the nontrivial $S=V/L$-wreath product, where $V=H\times U$. If $\mathcal{A}_{V}\neq\mathcal{A}_{H}\otimes\mathcal{A}_{U}$, then $\mathcal{A}_{V}$ is cyclotomic by Theorem 7.12(1). So Statement $(4)$ from Theorem 7.1 holds for $\mathcal{A}$ and $S$. Further, we assume that